# Is Parameter Collision Hindering Continual Learning in LLMs?

### v1

@@ Abstract
Shuo Yang1∗, Kun-Peng Ning1∗, Yu-Yang Liu1†, Jia-Yu Yao1, Yong-Hong Tian1, Yi-Bing Song2, Li Yuan1† 1School of Electronic and Computer Engineering, Peking University 2DAMO Academy, Alibaba Group {shuo_yang, ningkp}@stu.pku.edu.cn, {liuyuyang13, jiayu_yao, yhtian, yuanli-ece}@pku.edu.cn yibingsong.cv@gmail.com
These authors contributed equally to this work.Corresponding authors.
@@ 1 Introduction
To tackle this challenge in LLMs, existing studies Farajtabar et al. (2020); Saha et al. (2021); Wang et al. (2023a) attempted to employ orthogonal gradient descent-based methods to decouple parameter interdependence across different domains. Among them, O-LoRA (current SOTA CL method in LLMs) incrementally updates task-specific parameters by constraining them to the orthogonal subspace of previous tasks, ensuring that the LoRA parameters learned from past tasks remain fixed and independent across different subspaces Wang et al. (2023a). However, despite this independence, LoRA parameters within different subspaces may still interact, leading to potential parameter collisions.
In this paper, we reveal that constructing non-collision LoRA parameters provides a more effective solution to the problem of parameter interdependence, as non-collision serves as a sufficient but not necessary condition for orthogonality. As illustrated in Figure 1(a), assuming a three-dimensional parameter space, existing methods (e.g. O-LoRA) aim to learn three mutually orthogonal (blue box) task vectors $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$. Each task vector $[x_{i},y_{i},z_{i}]$ occupies all three subspaces of the parameters, and these parameters will interact with each other (red box), hurting the model performance. In contrast, as illustrated in Figure 1(b), non-collision constructs an extremely sparse parameter space, allowing each task vector to be updated in a local subspace without affecting each other. Non-collision parameters naturally achieve better task orthogonality (blue box), effectively decoupling parameter interdependence while preserving prior knowledge.
Leveraging the above insights, we propose Non-collision Low-Rank Adaptation (N-LoRA), a simple yet effective approach to tackling the continual learning problems in LLMs. Instead of directly optimizing the collision rate, we introduce $\ell_{1}$ constraints on task-specific LoRA parameters to make them extremely sparse, thereby significantly reducing the collision and dependency between various task vectors. As illustrated in Figure 1(c), compared with the orthogonal-based method (O-LoRA), the proposed non-collision-based approach (N-LoRA) can significantly achieve better task orthogonality (×$\bm{\times 4.1}$ times), lower parameter collision (×$\bm{\times 58.1}$ times), and superior average accuracy ($\bm{+2.9\%}$) in all cases.
We empirically validate the effectiveness of the proposed N-LoRA through extensive experiments, demonstrating that it significantly outperforms prior SOTA methods across various datasets and model architectures. The contribution of this paper can be summarized as follows:
@@
  • We introduce a simple yet effective method called N-LoRA for tackling the CL problem in LLMs, which leverages low collision parameters to enhance the model performance.
  • Our experiments on various CL benchmarks demonstrate that N-LoRA not only surpasses previous SOTA methods but also achieves better orthogonality and a lower collision rate.
@@ 2 Related Works
Parameter Efficient Fine-Tuning. Parameter Efficient Fine-Tuning (PEFT) He et al. (2021); Ding et al. (2022) enhances large language models (LLMs) performance by adjusting only a small portion of parameters, achieving results comparable to full fine-tuning while reducing computational costs. Key methods include Adapter tuning Houlsby et al. (2019), Prompt-based tuning Wei et al. (2021), P-tuning Liu et al. (2021a), Prefix tuning Li and Liang (2021), BitFit Zaken et al. (2021), and LoRA Wang et al. (2023a). In this work, we introduce a simple yet effective method built on LoRA to address this challenge, demonstrating strong performance.
Conventional Continual Learning. Conventional CL typically falls into three categories: (i) Rehearsal-based approaches Li and Hoiem (2017); Liu et al. (2021b); Lopez-Paz and Ranzato (2017) store and retrain on previous task data but may pose privacy concerns. (ii) Regularization-based approaches Kirkpatrick et al. (2017); Zenke et al. (2017) mitigate forgetting by imposing constraints that discourage significant changes in key weights. (iii) Architecture-based approaches Mallya and Lazebnik (2018); Wang et al. (2023b) expand model capacity dynamically or isolate existing weights with task-specific parameters to minimize interference between new and old tasks.
PEFT-Based Continual Learning for LLMs. LLMs with their vast number of parameters, create a significant computational burden for continual learning. To address this , methods like Learning to Prompt (L2P) Wang et al. (2022b) use a pool of small, learnable prompts to guide pre-trained models through tasks, while DualPrompt Wang et al. (2022a) combines expert and general prompts for task-specific and task-agnostic instructions. PEGP Qiao et al. (2024) utilizes Singular Value Decomposition(SVD) of task data to control gradient updates through orthogonal gradient projection, ensuring that new parameters do not interfere with the established feature space. O-LoRA Wang et al. (2023a) updates task-specific parameters by constraining them to an orthogonal subspace while keeping previous task parameters fixed However, O-LoRA struggles with parameter collisions, limiting its effectiveness in continual learning. Our proposed N-LoRA not only achieves better orthogonality than O-LoRA but also reduces parameter collisions and delivers superior performance.
@@ 3.1 Continual Learning for LLMs
Orthogonality-Based CL for LLMs. Continual learning in LLMs poses computational challenges due to their large parameter space. To mitigate this, existing studies have used PEFT methods and employed orthogonal gradient descent-based approaches to decouple parameter interdependence across different tasks. Among these, O-LoRA introduces matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ based on the low-rank assumption Wang et al. (2023a), updating the parameters as $W=W_{0}+\Delta W=W_{0}+AB$. O-LoRA enforces orthogonality between the task subspaces by imposing constraints on $A_{t}$ and $A_{i}$, introducing an additional orthogonality loss:
@@
We reveal that non-collision LoRA parameters provide a more effective solution to parameter interdependence and show that non-collision is a sufficient but not necessary condition for orthogonality (see Appendix 7.1 for proof). The detailed analysis is as follows,
@@ 3.3 Non-collision Low-Rank Adaptation
We attempt to sparsify individual LoRA parameters to minimize overall collisions. We show that the collision rate will decrease quadratically with the sparsity rate in Appendix 7.2. To further simplify, we introduce a sparsity constraint directly on $\Delta W_{i}$ while freezing the parameters from previous tasks. Specifically, for a sequence of tasks $T=\{t_{1},t_{2},\dots,t_{n}\}$, a corresponding sequence of task-specific matrices $\Delta W=\{\Delta W_{1},\Delta W_{2},\dots,\Delta W_{n}\}$ is obtained through continual fine-tuning using LoRA. For the $i$-th task, the matrix $\Delta W_{i}=A_{i}B_{i}$ is subject to $\ell_{1}$ regularization:
@@
Discussion: N-LoRA can effectively reduce the dimension of task subspaces. Beyond reducing collision rates, sparsity offers the additional advantage of effectively reducing the dimensionality of task subspaces. The nuclear norm, which provides an upper bound on the rank of a matrix Li and Chen (2021), corresponds to the number of subspaces utilized by each adapter in PEFT-based continual learning. Taking the first task of the Standard CL Benchmark as an example, we computed the nuclear norm of the LoRA weights inserted at each layer of the T5-large model for both O-LoRA and N-LoRA, as illustrated in Figure 3. With the rank set to 8 (i.e., 8 subspaces per task), O-LoRA (top plot) nearly fully occupied these subspaces, averaging 6.80 across all layers. In contrast, N-LoRA, by leveraging sparsity, utilized an average of only 0.14 subspaces. This clearly demonstrates that N-LoRA significantly reduces the dimensionality of task subspaces. Furthermore, knowledge from multiple domains will be preserved in these extremely low-dimensional (non-collision) parameter subspaces, making it more difficult to forget previously seen data.
@@ 4.1 Experimental Setups
Datasets. To evaluate our method, we use the Standard CL Benchmark Zhang et al. (2015) with five text classification datasets and the Large Number of Tasks Benchmark Razdaibiedina et al. (2023) encompassing 15 datasets, including GLUE Wang (2018), SuperGLUE Wang et al. (2019), and IMDB Maas et al. (2011). We follow the experimental setups of LFPT5 Qin and Joty (2021) and O-LoRA Wang et al. (2023a), using three different orders for each benchmark. Detailed task information and configurations are in Appendix 7.3, 7.4, and 7.5.
Metrics. Let $a_{i,j}$ be the test accuracy of the $i$-th task after training on the $j$-th task. $A_{i}$ denotes the LoRA $A$ matrix, and $\Delta W_{i}=A_{i}B_{i}$ matrix, both after training on the $i$-th task. We evaluate the model using the following six metrics:
@@
(\@slowromancapv@) Generalized Sparsity Rate (GSR) Hurley and Rickard (2009): This metric quantifies the sparsity of a matrix when none of its elements are zero. For detailed information and proof, see Appendix 7.6. The formula is:
@@ 4.2 Comparisons with SOTA methods
Performance on a Large Number of Tasks. Continual learning with a large number of tasks presents a significantly more challenging benchmark. As shown in Table 1, when the number of tasks increased, N-LoRA consistently outperforms all existing SOTA methods across all task orders (4, 5 and 6) in terms of average performance, with an average improvement of $\bm{2.8\%}$. However, despite this, N-LoRA, like other continual learning methods, still lags behind independent training approaches such as ProgPrompt and PerTaskFT, suggesting that there is considerable room for improvement in tackling a large number of tasks.

### v2

@@ Abstract
Shuo Yang1,∗, Kun-Peng Ning1,∗, Yu-Yang Liu1,†, Jia-Yu Yao1, Yong-Hong Tian1,2, Yibing Song3,4, Li Yuan1,2,† 1School of Electronic and Computer Engineering, Peking University 2Peng Cheng Laboratory 3DAMO Academy, Alibaba Group 4Hupan Lab {shuo_yang, ningkp}@stu.pku.edu.cn, {liuyuyang13, jiayu_yao, yhtian, yuanli-ece}@pku.edu.cn
These authors contributed equally to this work.Corresponding authors.The code and dataset are available at https://github.com/PKU-YuanGroup/N-LoRA.
@@ 1 Introduction
To tackle this challenge in LLMs, existing studies Farajtabar et al. (2020); Saha et al. (2021); Wang et al. (2023a) attempted to employ orthogonal gradient descent-based methods to decouple parameter interdependence across different domains. Among them, O-LoRA (the current SOTA CL method in LLMs) incrementally updates task-specific parameters by constraining them to the orthogonal subspace of previous tasks, ensuring the independence of LoRA parameters across subspacesWang et al. (2023a). However, despite this independence, interactions between LoRA parameters in different subspaces may still occur, potentially leading to parameter collisions.
In this paper, we reveal that constructing non-collision LoRA parameters provides a more effective solution to the problem of parameter interdependence, as non-collision serves as a sufficient but not necessary condition for orthogonality. By avoiding unnecessary overlaps in the parameter space, non-collision inherently minimizes interference among task-specific parameters, ensuring better task isolation and knowledge retention. As illustrated in Figure 1(a), assuming a three-dimensional parameter space, existing methods (e.g. O-LoRA) aim to learn three mutually orthogonal (blue box) task vectors $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$. Each task vector $[x_{i},y_{i},z_{i}]$ occupies all three subspaces of the parameters, and these parameters will interact with each other (red box), hurting the model performance. In contrast, as illustrated in Figure 1(b), non-collision constructs an extremely sparse parameter space, allowing each task vector to be updated in a local subspace without affecting each other. Non-collision parameters naturally achieve better task orthogonality (blue box), effectively decoupling parameter interdependence while preserving prior knowledge.
Leveraging the above insights, we propose Non-collision Low-Rank Adaptation (N-LoRA), a simple yet effective approach to tackling continual learning problems in LLMs. Instead of directly optimizing the collision rate, we introduce $\ell_{1}$ constraints on task-specific LoRA parameters to make them extremely sparse, thereby significantly reducing the collision and dependency between various task vectors. This sparsity ensures that task-specific parameters operate independently in dedicated subspaces, avoiding interference and allowing for seamless task transitions. As illustrated in Figure 1(c), compared to the orthogonal-based method (O-LoRA), the proposed non-collision-based approach (N-LoRA) can significantly achieve better task orthogonality (×$\bm{\times 4.1}$ times), lower parameter collision (×$\bm{\times 58.1}$ times), and superior average accuracy ($\bm{+2.9\%}$) in all cases.
We empirically validate the effectiveness of the proposed N-LoRA through extensive experiments, demonstrating that it significantly outperforms prior SOTA methods in various datasets and model architectures. The contribution of this paper can be summarized as follows:
@@
  • We introduce a simple yet effective method, called N-LoRA, to tackle the CL problem in LLMs, which leverages low collision parameters to enhance the model performance.
  • Our experiments on various CL benchmarks demonstrate that N-LoRA not only surpasses previous SOTA methods, but also achieves better orthogonality and a lower collision rate.
@@ 2 Related Works
Parameter Efficient Fine-Tuning. Parameter Efficient Fine-Tuning (PEFT) He et al. (2021); Ding et al. (2022) has become a vital area of research aimed at enhancing the performance of large language models (LLMs) by adjusting only a small portion of the model’s parameters. The primary objective of PEFT is to achieve performance on downstream tasks comparable to full model fine-tuning while minimizing computational resources by tuning only a small subset of model parameters. Various approaches have been proposed in this domain, including Adapter tuning Houlsby et al. (2019), Prompt-based tuning Wei et al. (2021), P-tuning Liu et al. (2021a), Prefix tuning Li and Liang (2021), BitFit Zaken et al. (2021), and LoRA Wang et al. (2023a). In this work, we introduce an easy yet effective method built on LoRA to address this challenge, demonstrating strong performance.
Conventional Continual Learning. Conventional CL typically falls into three categories: (i) Rehearsal-based approaches Li and Hoiem (2017); Liu et al. (2021b); Lopez-Paz and Ranzato (2017) store and retrain on previous task data, but can pose privacy concerns. (ii) Regularization-based approaches Kirkpatrick et al. (2017); Zenke et al. (2017) mitigate forgetting by imposing constraints that discourage significant changes in key weights. (iii) Architecture-based approaches Mallya and Lazebnik (2018); Wang et al. (2023b); Liu et al. (2021c, 2023); Ning et al. (2024) expand model capacity dynamically or isolate existing weights with task-specific parameters to minimize interference between new and old tasks.
PEFT-Based Continual Learning for LLMs. LLMs with their vast number of parameters, create a significant computational burden for continual learning. To address this, methods like Learning to Prompt (L2P) Wang et al. (2022b) use a pool of small, learnable prompts to guide pre-trained models through tasks, while DualPrompt Wang et al. (2022a) combines expert and general prompts for task-specific and task-agnostic instructions. PEGP Qiao et al. (2024) utilizes Singular Value Decomposition(SVD) of task data to control gradient updates through orthogonal gradient projection, ensuring that new parameters do not interfere with the established feature space. O-LoRA Wang et al. (2023a) incrementally updates new task parameters by constraining them to the orthogonal subspace of previous tasks while fixing the LoRA parameters learned from past tasks to minimize catastrophic forgetting. However, although O-LoRA attempts to update parameters within an orthogonal subspace, it struggles to prevent parameter collisions, which ultimately fails to resolve the forgetting problem. Our proposed N-LoRA not only achieves better orthogonality than O-LoRA but also reduces parameter collisions and delivers superior performance.
@@ 3.1 Continual Learning for LLMs
Orthogonality-Based CL for LLMs. Continual learning in LLMs poses computational challenges due to their large parameter space. To mitigate this, existing studies have used PEFT methods and employed orthogonal gradient descent-based approaches to decouple parameter interdependence across different tasks. Among these, O-LoRA introduces matrices $A\in\mathbb{R}^{d\times r}$ and, $B\in\mathbb{R}^{r\times k}$ based on the low-rank assumption Wang et al. (2023a), updating the parameters as $W=W_{0}+\Delta W=W_{0}+AB$. O-LoRA enforces orthogonality between the task subspaces by imposing constraints on $A_{t}$ and $A_{i}$, introducing an additional orthogonality loss:
@@
We reveal that non-collision LoRA parameters provide a more effective solution to parameter interdependence and show that non-collision is a sufficient but not necessary condition for orthogonality (see Appendix A.1 for proof). The detailed analysis is as follows,
@@ 3.3 Non-collision Low-Rank Adaptation
We attempt to sparsify individual LoRA parameters to minimize overall collisions. We show that the collision rate will decrease quadratically with the sparsity rate in Appendix A.2. To further simplify, we introduce a sparsity constraint directly on $\Delta W_{i}$ while freezing the parameters from previous tasks. Specifically, for a sequence of tasks $T=\{t_{1},t_{2},\dots,t_{n}\}$, a corresponding sequence of task-specific matrices $\Delta W=\{\Delta W_{1},\Delta W_{2},\dots,\Delta W_{n}\}$ is obtained through continual fine-tuning using LoRA. For the $i$-th task, the matrix $\Delta W_{i}=A_{i}B_{i}$ is subject to $\ell_{1}$ regularization:
@@
Discussion: N-LoRA can effectively reduce the dimension of task subspaces. Beyond reducing collision rates, sparsity offers the additional advantage of effectively reducing the dimensionality of task subspaces. The nuclear norm, which provides an upper bound on the rank of a matrix Li and Chen (2021), corresponds to the number of subspaces utilized by each adapter in PEFT-based continual learning. Taking the first task of the Standard CL Benchmark as an example, we computed the nuclear norm of the LoRA weights inserted at each layer of the T5-large model for both O-LoRA and N-LoRA, as illustrated in Figure 3. With the rank set to 8 (i.e., 8 subspaces per task), O-LoRA (top plot) nearly fully occupied these subspaces, averaging 6.80 across all layers. In contrast, N-LoRA, by leveraging sparsity, utilized an average of only 0.14 subspaces. This clearly demonstrates that N-LoRA significantly reduces the dimension of task subspaces. Furthermore, knowledge from multiple domains will be preserved in these extremely low-dimensional (non-collision) parameter subspaces, making it more difficult to forget previously seen data.
@@ 4.1 Experimental Setups
Datasets. To evaluate our method, we use the Standard CL Benchmark Zhang et al. (2015) with five text classification datasets and the Large Number of Tasks Benchmark Razdaibiedina et al. (2023) encompassing 15 datasets, including GLUE Wang (2018), SuperGLUE Wang et al. (2019), and IMDB Maas et al. (2011). We follow the experimental setups of LFPT5 Qin and Joty (2021) and O-LoRA Wang et al. (2023a), using three different orders for each benchmark. Detailed task information and configurations are in Appendix A.3, A.4, and A.6.
Metrics. Let $a_{i,j}$ be the test accuracy of the $i$-th task after training on the $j$-th task. $A_{i}$ denotes the LoRA $A$ matrix and $\Delta W_{i}=A_{i}B_{i}$ matrix, both after training on the $i$-th task. We evaluate the model using the following six metrics:
@@
(\@slowromancapv@) Generalized Sparsity Rate (GSR) Hurley and Rickard (2009): This metric quantifies the sparsity of a matrix when none of its elements are zero. For detailed information and proof, see Appendix A.7. The formula is:
@@
$ACR_{T}=\frac{2}{T(T-1)}\sum_{1\leq i<j\leq T}CR_{i,j}$ (10)
33footnotetext: With the exception of the results for O-LoRA and N-LoRA, which were obtained as part of this work, all other data in Table1 are based on those reported by Wang et al. (2023a), using the same experimental setup.
Implementation Details Our experiments on the T5-large model were conducted on machines equipped with NVIDIA A6000 GPUs. For tasks in orders 1 to 3, we used a learning rate of 1e-3, a batch size of 32, a dropout rate of 0.1, and a weight decay rate of 0. The models were trained for 10 epochs with $\lambda=0.4$. For tasks in order 4, we set the learning rate to 5e-4, the batch size to 32, the dropout rate to 0.1, and the weight decay rate to 0. The models were trained for 10 epochs, and $\lambda$ values were set as follows: (0.4, 0.4, 1.2, 0.4, 0.4, 0.4, 1.2, 1.2, 1.2, 0.4, 1.2, 1.2, 0.4, 0.4, 1.2). For tasks in order 5, we used a learning rate of 8e-4, a batch size of 32, a dropout rate of 0.1, and a weight decay rate of 0. The models were trained for 10 epochs with $\lambda=0.55$. For tasks in order 6, we set the learning rate to 1e-3, the batch size to 32, the dropout rate to 0.1, and the weight decay rate to 0. The models were trained for 10 epochs with $\lambda=1.8$.
Details for the specific task sequences in each order can be found in Appendix A.4. Details for experiments on the LLAMA-7B model can be found in Appendix A.6.
@@ 4.2 Comparisons with SOTA methods
Performance on a Large Number of Tasks. Continual learning with a large number of tasks presents a significantly more challenging benchmark. Among the evaluated methods, SeqLoRA exhibits the worst performance due to significant parameter collisions and overlaps in non-orthogonal spaces. Its metrics, including OO (3796.23), AWOM (1120.5), Sparsity Rate (0.682), and ACR (0.542), reflect severe inefficiencies, limiting its ability to maintain task-specific knowledge.
In contrast, as shown in Table 1, N-LoRA surpasses all existing SOTA methods across task orders (4, 5, and 6), achieving an average performance improvement of $\bm{2.8\%}$. By reducing parameter collisions and optimizing subspace allocation, N-LoRA delivers significantly better metrics, outperforming SeqLoRA by factors of approximately 1k, 8k, 85, and 71, respectively. These results underscore N-LoRA’s superior capacity to handle numerous tasks while preserving task-specific knowledge. Further comparisons with O-LoRA are detailed in Appendix A.5.